# Guangyu Zhu, Yiqin Deng, , Xianhao Chen, , Haixia Zhang, , Yuguang Fang, , Tan F. Wong

### v1

@@ IV Optimization and Solution Approach
The ultimate objective of the ESFL training is to minimize the total training time in each round given by (as we mentioned earlier, we will omit the dependence of the round number $r$ for notational simplicity)
$\begin{gathered}\min T=\min\max_{i}T_{i}.\end{gathered}$ (9)
@@
Then, our ESFL optimization problem can be transformed into a mixed-integer non-linear program (MINLP), which is typically NP-hard. To solve the problem efficiently, we decompose it into two subproblems and solve them iteratively. We construct the first subproblem for cut layer decision by treating computing resource allocation as variables:
$\begin{gathered}\min_{C_{i}}T^{e}_{i}=\frac{w_{c}^{l_{i}}\cdot n_{i}}{c_{i}}+\frac{w_{b}^{l_{i}}\cdot n_{i}}{b^{u}_{i}}+\\ \frac{(D-w_{c}^{l_{i}})\cdot n_{i}}{C_{i}}+\frac{w_{b}^{l_{i}}\cdot n_{i}}{b^{d}_{i}}\\ s.t.~{}\sum_{i=0}^{S}C_{i}\leq C_{total},\end{gathered}$ (12)
where communication and computing workload for all users are constants since the cut layers $l_{i}$ are predetermined by solving the previous subproblem.
The alternative optimization algorithm is shown in Algorithm 3. For the notational convenience, we use $\{C_{*}\}$ to denote $\{C_{1},C_{2},...,C_{S}\}$ and $\{l_{*}\}$ denote $\{l_{1},l_{2},...,l_{S}\}$.
@@ V-A Experimental Setup
For all following experiments, we evaluate the performance on image classification tasks over the common dataset CIFAR-10, and leverage VGG13, VGG16 and VGG19 [25] framework as the neural network architecture to implement the distributed applications. We compare our ESFL with the alternatives such as FedAVG (FL), original split learning (SL), splitfed learning (SFL).
Dataset CIFAR-10 [26] contains $50,000$ training samples and $10,000$ testing samples in $10$ classes, with $6,000$ images per class. We assume that there are $100$ users participating in the whole training, while only $10$ users will be randomly selected to join one-round training. Under the assumption that all users’ data samples are independently and identically distributed (IID), the data is shuffled and then partitioned into $100$ clients, each receiving $500$ samples.
Training Configuration We use a distributed machine learning framework, similar to federated learning, which has several learning hyperparameters including local learning rate $\rho_{r}$, where $\rho_{0}=0.01$ and $\rho_{r}$ will be decaying as the round number $r$ increases, and the constant number of local epochs $\epsilon_{i}=5$. Moreover, we introduce a global learning rate $\eta=0.5$ to control the global model updating pace. According to the experimental results, when choosing a mini-batch size of $32$, we can obtain a well trained model.
Neural Network Architecture We deploy the VGG19 network [25] as the training model, which primarily consists of convolutional layers (CONV), fully-connected layers (FC), and softmax layer (SoftMax). We resize the input layer of the original VGG13, VGG16 and VGG19 from $224\times 224$ to $32\times 32$ to fit the CIFAR-10 dataset. The mini-batch size is set to 32. We present VGG19 architecture and workload of each layer in Table I.
TABLE I: VGG19 Network Architecture and Parameters
@@ V-B Model Performance
The key hyperparameter in our ESFL that affect the final convergence performance such as testing accuracy and loss is the number of training rounds, since we leverage FedAVG for all distributed ML algorithms except SL. For a fair comparison, we set the training threshold for VGG13 to $88\%$, VGG16 to $87.5\%$ and VGG19 to $86.5\%$ testing accuracy based on the worst converged accuracy. Three distributed ML algorithms (FL, SFL, ESFL) achieve the expected convergence performance at the $1500$-th training round, and SL achieves the expected convergence performance at the $200$-th training round. The testing performance results are shown in Figure. 4.
@@ V-C1 Resource Limitation
We separate the impacts of user-side communication and computing resource limitation by simulating four resource settings shown in Table II: Both Poor (BP) indicates that both communication and computing resources are highly limited at EDs, Poorcom Richcmp (PR) indicates that communication resources are highly limited while computing resources are slightly limited (five times larger than that for the highly limited case), Richcom Poorcmp (RP) indicates that communication resources are slightly limited while computing resources are highly limited, and Both Rich (BR) indicates that both communication and computing resources are slightly limited. The selection process for each user-side resource setting is similar to Table II. For learning algorithmic implementation, we use original FL and SL, and SFL, which is similar to that for ESFL introduced in Section III.
TABLE II: Communication and computing resource settings
@@
Table III presents the average one-round training and communication time and one-round communication time, for different NNs under different resource scenarios using FL, SL, SFL and efficient split federated learning ESFL, respectively. Figure. 5 shows the allocation results of user-side training workload represented as cutting-layer distributions. From cutting-layer distributions, as the user-side resource becomes more sufficient, our ESFL applies more identical cutting-layer distributions strategies for all NNs. For the results in those two table, our ESFL algorithm significantly reduces one-round training and communication time under all circumstances. These advantages often stem from the dynamics between user-side computing and communication resource. In scenarios where local computing resource are poor whereas communication resource are rich (RP), ESFL remains less layers of user-side models leveraging more on server-side computing power. Conversely, under the PR scenario, ESFL mitigates these limitations by remaining more layers of user-side model to rely less on communication. Comparing one-round training and communication time of SFL and FL, an intriguing phenomenon emerges. Although SFL leverages the server-side resource to accelerate training, improper user-side workload allocation (model separation) and server-side resource allocation lead to decreased time efficiency.
TABLE III: One-round training and communication time for different resource settings
@@
Figure 5: Cutting-layer distributions (user-side workloads allocation) of three NNs under four different resource limitations using ESFL algorithm.
Table IV indicates the total training and communication time for four ML algorithms to achieve the expected convergence performance. In the context of contrasting FL and SFL, it is imperative to acknowledge that although SFL, similar to ESFL, utilizes server-side computing resource, its overall performance is significantly influenced by by the harmonization of user-side workload and server-side resource allocation. The lack of effective resource allocation strategies can result in inferior performance in SFL when compared to FL. Especially in RP and BR, when communication resource are notably limited in EDs, FL outperforms SFL. This performance discrepancy can be attributed to the lack of effective allocating strategies, which results in SFL demonstrating inferior performance when compared to FL. Nevertheless, our ESFL exhibits a significant increase in efficiency compared to both FL and SFL across all tested scenarios. This provides the evidence that a well-conceived strategy for workload and resource allocation can markedly enhance the efficiency of the whole training process. While the model used here and other ML algorithms are not the state-of-the-art for this task, it does provide sufficient evidence to show that our ESFL can significantly reduce training latency and improve training efficiency by considering user-side resource heterogeneity.
TABLE IV: Total training and communication time for different resource settings
@@ V-C2 Resource Heterogeneity
In this section, we assume there exist four different communication and computing resource settings to evaluate the efficiency of our ESFL algorithm under different heterogeneous scenarios. The detail of our resource simulation setting is shown in Table V, where in every training round, the available communication and computing conditions of the selected users are randomly chosen from resource options in one scenario. In particular, considering the fairness, the average resource amounts in different scenarios are equal, and only resource distributions are dissimilar to simulate different resource heterogeneity. Four heterogeneous scenarios are considered: Small Heterogeneity (SH), implying that both communication and computing resource heterogeneity at EDs is small, Smallcom Largecmp (SL), implying that communication heterogeneity is small while computing resource is large, Largecom Smallcmp (LS), implying that communication heterogeneity is large while computing resource is small, and Large Heterogeneity (LH), implying that both communication and computing resource heterogeneities are large. For instance, in SH, each user will randomly choose one communication condition from [$10$, $15$, $20$, $25$] kiloBytes (KBps) and one computing condition from [$1.3$, $1.95$, $2.6$, $3.25$] TFLOPs as their available resources, and the server will base on the resource information of the selected users to allocate appropriate server-side computing resource and make user-side cutting layer decision for all selected users. To simulate the training workload heterogeneity, we assume all available users have heterogeneous but constant amounts of data samples, which are chosen from [$200$, $400$, $600$, $800$].
TABLE V: Communication and computing resource settings
@@
Table VI presents the average one-round training and communication time and one-round communication time, for different NNs under different resource scenarios using FL, SL, SFL and ESFL, respectively. For the results in those two table, our ESFL algorithm significantly reduces one-round training and communication time under all circumstances and is least affected by resource heterogeneity, where both communication and computing heterogeneity seriously impact the training efficiencies of the other three ML algorithms. It is noteworthy that as the distribution of resources approaches a state of greater uniformity (SH), the gap of time efficiency between SFL and ESFL is decreased. Conversely, with an increase in resource heterogeneity (LH), the performance differential between ESFL and SFL widens notably, which highlight the robustness of ESFL in diverse resource environments. For instance, the training latency of SFL in LH is increased by nearly two times compared with that in SH while training latencies of ESFL are nearly the same in all scenarios. The user-side training workload allocation results (cutting-layer distributions) is shown in Figure. 6. ESFL algorithm demonstrates a trend implementing more uniform cutting-layer distributions across all NNs, under more identical resource distributions.
TABLE VI: One-round training and communication time for different heterogeneities
@@
Figure 6: Cutting-layer distributions (user-side workloads allocation) of three NNs under four different resource heterogeneities using ESFL algorithm.
Table VII indicates the total training and communication time under different RH. It can be seen from the results that our ESFL method is significantly more efficient compared with the original FL, SL and SFL in most scenarios, only except VGG13 in SL. Comparing communication time of FL in SL and LS, the performance of FL is markedly impacted by the RH. However, ESFL capitalizes on these heterogeneities through strategic allocation: in the environment with low communication heterogeneity (SL), it increases the user-side communication workload while decreases computing workload; in the environment with high communication heterogeneity (LS), it conversely allocates more computing workload to EDs. Therefore, our approach optimizes the time efficiency across diverse scenarios by adaptively leveraging RH.
TABLE VII: Total training time to achieve the convergence performance for different heterogeneities
@@ V-D Resource Allocation Convergence Analysis
From Section IV, the joint resource allocation and model splitting problem has been decomposed into two subproblems due to the time complexity, and we leverage an alternative optimization approach to solve those subproblems. To evaluate the convergence of our alternative method, we simulate iterative results at four users’ scales, 100 users, 200 users, 400 users and 800 users. The simulation result in Figure. 7 shows that even for the largest user scale (800 users), our alternative approach only needs $10$ iterations to achieve convergence, while when the user scales are small, it only requires a few iterations to achieve convergence.
Figure 7: Average one-round training time using iterative optimization.

### v2

@@ IV Optimization and Solution Approach
The ultimate objective of the ESFL training is to minimize the total training time. However, since every round training time are independent, minimizing the total training time is equivalent to minimizing the total training time in each round including computing and communication time given by (as we mentioned earlier, we will omit the dependence of the round number $r$ for notational simplicity)
$\begin{gathered}\min T=\min\sum_{r}\max_{i}T_{i}.\end{gathered}$ (9)
@@
The alternative optimization algorithm is shown in Algorithm 3. For the notational convenience, we use $\{C_{*}\}$ to denote $\{C_{1},C_{2},...,C_{S}\}$ and $\{l_{*}\}$ denote $\{l_{1},l_{2},...,l_{S}\}$. The reason why we use an alternative optimization approach is that the cut layer decisions for different selected users lead to the variance at the user-side workload. Moreover, the allocation of server-side computing resource to individual one user will affect the availability of resources for others, given the fixed total capacity of server-side resources, where the joint optimization of worklaods and resources incurs the coupling effect for different users. To address this coupling effect, our ESFL algorithm transforms the optimization problem into a mixed-integer non-linear program (MINLP), which is typically NP-hard. To solve the problem efficiently, we decompose it into two subproblems and solve them iteratively. We construct the first subproblem for cut layer decision by treating computing resource allocation as fixed decision variables:
$\begin{gathered}\min_{C_{i}}\max T^{e}_{i}=\frac{w_{c}^{l_{i}}\cdot n_{i}}{c_{i}}+\frac{w_{b}^{l_{i}}\cdot n_{i}}{b^{u}_{i}}+\\ \frac{(D-w_{c}^{l_{i}})\cdot n_{i}}{C_{i}}+\frac{w_{b}^{l_{i}}\cdot n_{i}}{b^{d}_{i}}\\ s.t.~{}\sum_{i=0}^{S}0\leq C_{i}\leq C_{total},\end{gathered}$ (12)
where communication and computing workloads for all users are constant since the cut layers $l_{i}$ are predetermined by solving the previous subproblem. Plus, the downlink $b^{d}_{i}$, uplink communication resource $b^{u}_{i}$ and user-side available computing resource $c_{i}$ are constant. Therefore, the equation (12) can be abbreviated as:
$\begin{gathered}\min_{C_{i}}\max T^{e}_{i}=\frac{a}{C_{i}}+b,\end{gathered}$ (13)
where $a$ and $b$ are constant. To solve this min-max problem, we assume there exists a variable $K$, where for all $C_{i}$, $K\geq\frac{a}{C_{i}}+b$. Then, we construct the equation(13) as a minimizing problem:
$\begin{gathered}\min K\\ s.t.~{}\sum_{i=0}^{S}0\leq C_{i}\leq C_{total},\\ K\geq\frac{a_{1}}{C_{1}}+b_{1}\\ K\geq\frac{a_{2}}{C_{2}}+b_{2}\\ \vdots\\ K\geq\frac{a_{|S|}}{C_{|S|}}+b_{|S|}.\end{gathered}$ (14)
When $C_{i}\geq 0$, $T^{e}_{i}$ is a convex function ($\nabla^{2}T^{e}_{i}\geq 0$). Then, a convex optimization solver [29] can be leveraged to solve this subproblem.
@@ V-A Experimental Setup
For all following experiments, we evaluate the performance on image classification tasks over the common dataset CIFAR-10 and leverage VGG13, VGG16 and VGG19 [30] framework as the neural network architecture to implement the distributed applications. We compare our ESFL with the alternatives such as FedAVG (FL), original split learning (SL), and splitfed learning (SFL).
Dataset: CIFAR-10 [31][32] contains $50,000$ color training images and $10,000$ testing images with $32\times 32$ resolution in $10$ classes, with $6,000$ images per class. We assume that there are $100$ users participating in the whole training process, while only $10$ users are randomly selected to join one-round training. Under the assumption that all users’ data samples are independently and identically distributed (IID), the data is shuffled and then partitioned into $100$ clients with no replacement, every user owning $500$ training samples.
Training Configuration: We use a distributed machine learning framework, similar to federated learning, which has several learning hyperparameters including local learning rate $\rho_{r}$, where $\rho_{0}=0.01$ and $\rho_{r}$ is decaying as the round number $r$ increases and the constant number of local epochs $\epsilon_{i}=5$. Moreover, we introduce a global learning rate $\eta=0.5$ to control the global model updating pace. According to the experimental results, when choosing a mini-batch size of $32$, we can obtain a well trained model.
Neural Network Architecture: We deploy the VGG19 network [30] as the training model, which primarily consists of convolutional layers (CONV), fully-connected layers (FC), and softmax layer (SoftMax). We resize the input layer of the original VGG13, VGG16, and VGG19 from $224\times 224$ to $32\times 32$ to fit the CIFAR-10 dataset. The mini-batch size is set to 32. We present VGG19 architecture and workload of each layer in Table II.
TABLE II: VGG19 Network Architecture and Parameters
@@ V-B Model Performance
One key hyperparameter in our ESFL that affect the final convergence performance such as testing accuracy and loss is the number of training rounds, since we leverage FedAVG for all distributed ML algorithms except SL. For a fair comparison, we set the training threshold for VGG13 to $88\%$, VGG16 to $87.5\%$ and VGG19 to $86.5\%$ testing accuracy based on the worst converged accuracy. Three distributed ML algorithms (FL, SFL, ESFL) achieve the expected convergence performance at the $1500$-th training round, and SL achieves the expected convergence performance at the $200$-th training round. The testing performance results are shown in Figure. 4.
@@ V-C1 Resource Limitation
TABLE III: Communication and computing resource settings
@@
TABLE IV: One-round training and communication time for different resource settings
@@
We separate the impacts of user-side communication and computing resource limitation by simulating four resource settings shown in Table III: Both Poor (BP) indicates that both communication and computing resources are highly limited at EDs, Poorcom Richcmp (PR) indicates that communication resources are highly limited while computing resources are slightly limited (five times larger than that for the highly limited case), Richcom Poorcmp (RP) indicates that communication resources are slightly limited while computing resources are highly limited, and Both Rich (BR) indicates that both communication and computing resources are slightly limited. The selection process for each user-side resource setting is similar to Table III. For learning algorithmic implementation, we use original FL and SL, and SFL, which is similar to that for ESFL introduced in Section III.
@@
Figure 5: Cut layer distributions (user-side workloads allocation) of three NNs under four different resource limitations using ESFL algorithm.
Table IV presents the average one-round training and communication time and one-round communication time for different NNs under different resource scenarios using FL, SL, SFL and efficient split federated learning ESFL, respectively. Figure. 5 shows the allocation results of user-side training workload represented as cut layer distributions. The cut layer distribution represents the empirical probability of selecting layer $l$ for user $i$ in the total training rounds, which is $P_{i,l}=\sum^{R}_{r}\frac{x_{i,r}^{l}}{R}$, where $x_{i,r}^{l}$ is the cutting layer decision showing in equation 8 and $\sum_{l}P_{i,l}=1$. The cut layer distribution combining with the amounts of user-side data indicates the allocated user-side computing and communication workload. Therefore, from cut layer distributions, as the user-side resource becomes more sufficient, our ESFL applies more identical cut layer distributions strategies for all NNs. For the results in those two tables, our ESFL algorithm significantly reduces one-round training and communication time under all circumstances. These advantages often stem from the dynamics between user-side computing and communication resource. In scenarios where local computing resource are poor whereas communication resource are rich (RP), ESFL remains fewer layers of user-side models by leveraging more on server-side computing power. Conversely, under the PR scenario, ESFL mitigates these limitations by remaining more layers of user-side model to rely less on communication. Comparing one-round training and communication time of SFL and FL, an intriguing phenomenon emerges. Although SFL leverages the server-side resource to accelerate training, improper user-side workload allocation (model separation) and server-side resource allocation lead to decreased time efficiency.
TABLE V: Total training and communication time for different resource settings
@@
Table V indicates the total training and communication time for four ML algorithms to achieve the expected convergence performance. In the context of contrasting FL and SFL, it is imperative to acknowledge that although SFL, similar to ESFL, utilizes server-side computing resource, its overall performance is significantly influenced by the harmonization of user-side workload and server-side resource allocation. The lack of effective resource allocation strategies can result in inferior performance in SFL when compared to FL. Especially in RP and BR, when communication resource is notably limited in EDs, FL outperforms SFL. This performance discrepancy is attributed to the lack of effective allocating strategies, which results in inferior performance in SFL when compared to FL. Nevertheless, our ESFL exhibits a significant increase in efficiency compared to both FL and SFL across all tested scenarios. This provides the evidence that a well-conceived strategy for workload and resource allocation can markedly enhance the efficiency of the whole training process. While the model used here and other ML algorithms are not the state-of-the-art for this task, it does provide sufficient evidence to show that our ESFL can significantly reduce training latency and improve training efficiency by considering user-side resource heterogeneity.
@@ V-C2 Resource Heterogeneity
TABLE VI: Communication and computing resource settings
@@
TABLE VII: One-round training and communication time for different heterogeneities
@@
Figure 6: Cut layer distributions (user-side workloads allocation) of three NNs under four different resource heterogeneities using ESFL algorithm.
In this section, we assume there exist four different communication and computing resource settings to evaluate the efficiency of our ESFL algorithm under different heterogeneous scenarios. The detail of our resource simulation setting is shown in Table VI, where in every training round, the available communication and computing conditions of the selected users are randomly chosen from resource options in one scenario. In particular, considering the fairness, the average resource amounts in different scenarios are equal, and only resource distributions are dissimilar to simulate different resource heterogeneity. Four heterogeneous scenarios are considered: Small Heterogeneity (SH), implying that both communication and computing resource heterogeneity at EDs is small, Smallcom Largecmp (SL), implying that communication heterogeneity is small while computing resource is large, Largecom Smallcmp (LS), implying that communication heterogeneity is large while computing resource is small, and Large Heterogeneity (LH), implying that both communication and computing resource heterogeneities are large. For instance, in SH, each user will randomly choose one communication condition from [$10$, $15$, $20$, $25$] kiloBytes (KBps) and one computing condition from [$1.3$, $1.95$, $2.6$, $3.25$] TFLOPs as their available resources, and the server will base on the resource information of the selected users to allocate appropriate server-side computing resource and make user-side cutting layer decision for all selected users. To simulate the training workload heterogeneity, we assume that all available users have heterogeneous but constant amounts of data samples, which are chosen from [$200$, $400$, $600$, $800$].
Table VII presents the average one-round training and communication time and one-round communication time, for different NNs under different resource scenarios using FL, SL, SFL and ESFL, respectively. For the results in Table VII, our ESFL algorithm significantly reduces one-round training and communication time under all circumstances and is least affected by resource heterogeneity, where both communication and computing heterogeneity seriously impact the training efficiencies of the other three ML algorithms. It is noteworthy that as the distribution of resources approaches a state of greater uniformity (SH), the gap of time efficiency between SFL and ESFL is decreased. Conversely, with an increase in resource heterogeneity (LH), the performance differential between ESFL and SFL widens notably, which highlights the robustness of ESFL in diverse resource environments. For instance, the training latency of SFL in LH is increased by nearly two times compared with that in SH while training latencies of ESFL are nearly the same in all scenarios. The user-side training workload allocation results (cut layer distributions) is shown in Figure. 6. From this simulation result, ESFL algorithm demonstrates a trend implementing more uniform cut layer distributions across all NNs, under more identical resource distributions.
TABLE VIII: Total training time to achieve the convergence performance for different heterogeneities
@@
Table VIII presents the total training and communication time under different RH. It can be seen from the results that our ESFL method is significantly more efficient compared with the original FL, SL, and SFL in most scenarios, only except VGG13 in SL. Comparing communication time of FL in SL and LS, the performance of FL is markedly impacted by the RH. However, ESFL capitalizes on these heterogeneities through joint workload and resource allocation: in the environment with low communication heterogeneity (SL), it increases the user-side communication workload while decreases computing workload; in the environment with high communication heterogeneity (LS), it conversely allocates more computing workload to EDs. Therefore, our approach optimizes the time efficiency across diverse scenarios by adaptively leveraging RH.
@@ V-D Resource Allocation Convergence Analysis
(a) Average one-round training time in BP
(b) Average one-round training time in PR
(c) Average one-round training time in RP
(d) Average one-round training time in BR
Figure 7: Average one-round training time using iterative optimization of VGG19 in four different resource scenarios shown in Table III.
From Section IV, the joint resource allocation and model splitting problem has been decomposed into two subproblems due to the time complexity, and we leverage an alternative optimization approach to solve those subproblems. To evaluate the convergence of our alternative method, we simulate iterative results at four users’ scales in four resource scenarios (Table III), 100 users, 200 users, 400 users, and 800 users. The simulation results in Figure. 7 shows that even for the largest user scale (800 users) in all different resource scenarios, our alternative approach only needs $9$ iterations to achieve convergence, while when the user scales are small, it only requires a few iterations to achieve convergence. Comparing average one-round training time across different scenarios, our method significantly enhances training efficiency, particularly in the resource-constrained scenario (BP).